OpenStudy (anonymous):

harry bought 6 hotdogs and 4 sodas and paid 6.70. jassi bought 4 hotdogs and 3 drinks and paid 4.65

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

6 hotdogs and 4 drinks for 6.70 4 hotdogs and 3 drinks for 4.65

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So where is the question?

OpenStudy (anonymous):

right there above

OpenStudy (anonymous):

What do you expect me to do with it?

OpenStudy (anonymous):

help me solve it please its systemof equations

OpenStudy (anonymous):

I know, what do you want me to solve for?

OpenStudy (anonymous):

the price of each

OpenStudy (anonymous):

Is this online, or on paper?

OpenStudy (anonymous):

paper

OpenStudy (anonymous):

Oh, okay.

OpenStudy (anonymous):

Give me a sec.

OpenStudy (anonymous):

Substitution?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I will teach you another method. I don't know what it is called, but it works here.

OpenStudy (anonymous):

Set it up with drinks being y and food being x.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

And set it equal to the total.

OpenStudy (anonymous):

2 equations.

OpenStudy (anonymous):

1 on top of the other.

OpenStudy (anonymous):

yea thats how i have it

OpenStudy (anonymous):

i just dont know how to solve it

OpenStudy (anonymous):

Okay, subtract 1 from the other. Each term

OpenStudy (anonymous):

You should get 2x+y=2.05

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

wait subtrct a dollar

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

i got x=19.35 and y=24.25 is that right

OpenStudy (anonymous):

You added, you need to subtract.

OpenStudy (anonymous):

Wait, how did you get there?

OpenStudy (anonymous):

subtract which one

OpenStudy (anonymous):

Can't skip, sorry.

OpenStudy (anonymous):

(I can't)

OpenStudy (anonymous):

what do u mean cant skip

OpenStudy (anonymous):

I think you skipped some steps.

OpenStudy (anonymous):

oh ok 6x+4y=6.70 first problem right

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Then solve for y in the 2x+y=2.05

OpenStudy (anonymous):

where did you get that 2x+y=2.05

OpenStudy (anonymous):

Subtract both equations

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

Then solve for y.

OpenStudy (anonymous):

Then substitute.

OpenStudy (anonymous):

Rob, I'm doing this correctly, right?

OpenStudy (anonymous):

Solve {6 hotdogs + 4 drinks == 6.70, 4 hotdogs + 3 drinks == 4.65} for hotdogs and drinks {hotdogs = 0.75, drinks = 0.55}

OpenStudy (anonymous):

do i divide 2x by 2.05

OpenStudy (anonymous):

to get y

OpenStudy (anonymous):

No, subtract 2x.

OpenStudy (anonymous):

I need to finish HW, listen to Rob.

OpenStudy (anonymous):

Bye.

OpenStudy (anonymous):

y=2x+2.05 right rob